OpenStudy (jh99):

There is no way I can find the intercept form of y=-2x^2-4x+3 ... What am I supposed to do?

satellite73 (satellite73):

what does "intercept form" mean for a quadratic?

satellite73 (satellite73):

does it mean "factored form"?

OpenStudy (mathmate):

Intercept form : y=a(x-x1)(x-x2)

OpenStudy (jh99):

I'm sure they're asking for the factored form: y=a(x+r1)(x+r2)

OpenStudy (jh99):

sorry minus*

OpenStudy (mathmate):

So @satellite73 's suggested "factored form" is the same.

satellite73 (satellite73):

does not factor using integers

OpenStudy (jh99):

so it's basically impossible?

satellite73 (satellite73):

using integers, yes

OpenStudy (mathmate):

@jh99 as @satellite73 said, it is impossible to express the given equation y=\(-2x^2-4x+3\) in intercept form using rational number. However, either you find a typo in the question, or you'd have to accept the irrational version of the intercept form, which is not impossible: You can factorize by using the quadratic formula, which gives \(x=\{-(\sqrt{10}+2)/2, (\sqrt{10}-2)/2\}\) which gives \(y=-\frac{1}{2}(2x+\sqrt{10}+2)(2x-\sqrt{10}+2)\)